OpenStudy (anonymous):

Evaluate: X ÷ YZ, for X=3 3/8, Y = -6, and Z = 2 9/10

OpenStudy (anonymous):

I got -261/160. is this right

OpenStudy (anonymous):

Transform the mixed fraction into improper and then multiply by the reciprocal to get the fraction out of the denominator

OpenStudy (anonymous):

that is not correct \[\frac{3\frac{3}{8}}{-6*2\frac{9}{10}}\] this is the first step then place into improper fractions

OpenStudy (anonymous):

I did those steps

OpenStudy (anonymous):

27/8 ÷ -6/1

OpenStudy (anonymous):

\[\frac{\frac{27}{8}}{-6*\frac{29}{10}}\] the top part is correct but no the bottom fraction look at the denominator above